# On the Source Model Key Agreement Problem

### v2

@@ Appendix C An equivalent form of Conjecture 1
Note that by setting $Z=Z^{\prime}$, $Y=Y^{\prime}$ and replacing $X$ by $(X,X^{\prime})$ in Conjecture 1, we obtain the following equation (as a special case):
$S(XX^{\prime};Y^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}),\qquad\forall P_{XX^{\prime}Y^{\prime}Z^{\prime}}.$ (92)
Observe that $S(XX^{\prime};Y^{\prime}\|Z^{\prime})\geq S(XX^{\prime};XY^{\prime}\|XZ^{\prime})$ because the first terminal can publicly reveal $X$ to all parties. Therefore, (92) can be weakened to
$S(XX^{\prime};XY^{\prime}\|XZ^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}),\qquad\forall P_{XX^{\prime}Y^{\prime}Z^{\prime}}.$ (93)
@@
$S(X^{\prime};YY^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X^{\prime}Z^{\prime};Y|Y^{\prime}).$ (94)
@@
$S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(XX^{\prime};Y^{\prime}|Z^{\prime})\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime}).$ (95)
Adding up (92) and (95) we obtain
$S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (96)
@@
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}|Z^{\prime})$ $=S(X;Y\|Z)-S(XX^{\prime};YY^{\prime}|Z^{\prime})+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S(X;Y\|Z)-S(X;Y|Z^{\prime})+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$ (97)
$\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime})$ (98)
@@
$\frac{1}{n}H(K_{B})$ $=\frac{1}{n}I(X^{n}F_{1:r}W_{1};K_{B})+\epsilon_{n}$ (100)
$=\frac{1}{n}(I(X^{n}F_{1:r}W_{1};K_{B})-I(K_{B};Z^{n}F_{1:r}))+2\epsilon_{n}$ (101)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r}W_{1};K_{B}\|Z^{n}F_{1:r})+2\epsilon_{n}$ (102)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r}W_{1};Y^{n}F_{1:r}W_{2}\|Z^{n}F_{1:r})+2\epsilon_{n}$ (103)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r-1}W_{1};Y^{n}F_{1:r-1}W_{2}\|Z^{n}F_{1:r-1})+2\epsilon_{n}$ (104)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r-2}W_{1};Y^{n}F_{1:r-2}W_{2}\|Z^{n}F_{1:r-2})+2\epsilon_{n}$ (105)
[…]
$\leq\frac{1}{n}\Phi(X^{n}W_{1};Y^{n}W_{2}\|Z^{n})+2\epsilon_{n}$ (107)
@@
where $\epsilon_{n}$ converges to zero as $\delta$ converges to zero, (100) follows from the fact that $K_{A}=K_{B}$ with probability $1-\delta$ and $H(K_{A}|X^{n}F_{1:r}W_{1})=0$, (101) follows from the privacy constraint, (102) from property iv in Lemma 4, (103) follows from property iii in Lemma 4 and the fact that $K_{B}$ is a function of $(Y^{n}W_{2}F_{1:r})$, (104)-(107) follow from property ii in Lemma 4 and the fact that $H(F_{i}|X^{n},F^{i-1},W_{1})=0$ for odd $i$ and $H(F_{i}|Y^{n},F^{i-1},W_{2})=0$ for even $i$, (108) follows from property v in Lemma 4, and (109) follows from property i in Lemma 4.
@@
Observe that $S(X;Y\|Z)$ similar properties as above properties, i.e., $S(X^{n};Y^{n}\|Z^{n})=nS(X;Y\|Z)$, and $S(F,X;F,Y\|F,Z)\leq S(F,X;Y\|Z)$, etc; see [14].
@@
Take $n$, i.i.d copies of $(X,Y,Z,\tilde{X})$ and denote it by $(X^{n},Y^{n},Z^{n},\tilde{X}^{n})$. We have:
@@
The Markov chain $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ implies the Markov chain $(F,\tilde{X})-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$. Let $X^{\prime}=(F,\tilde{X})$. We have
@@
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;F,Y,Z|F,\tilde{X})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;Y,Z|\tilde{X})-I(F;YZ|\tilde{X})$
$\leq S(\tilde{X};Y\|Z)+I(F,X;Y,Z|\tilde{X})$ (111)
@@
From $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$, we deduce $(F,\tilde{X})-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$ holds. Let $X^{\prime}=(F,\tilde{X})$. Thus, we have:
@@
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;F,Y,Z|F,\tilde{X})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(X;F,Y,Z|\tilde{X})-I(X;F|\tilde{X})$
$\leq S(\tilde{X};F,Y\|Z)+I(X;F,Y,Z|\tilde{X})$
@@
Proof of Property iv) We know that $S(X^{\prime};Y\|Z)\geq I(X^{\prime};Y)-I(Y;Z)$. Thus $S(X^{\prime};Y\|Z)+I(X;Y,Z|X^{\prime})\geq I(X^{\prime};Y)-I(Y;Z)+I(X;Y,Z|X^{\prime})=I(X;Y)-I(Y;Z)+I(X;Z|X^{\prime},Y)\geq I(X;Y)-I(Y;Z).$ This completes the proof.
Proof of Property v) Take some arbitrary $P_{X^{\prime}|X}$ reaching within $\epsilon$ of the infinimum defining $\Phi(X;Y|Z)$. If we define $(W_{1},W_{2})$ independent of $(X^{\prime},X,Y,Z)$, the Markov chain $X^{\prime}-\!\!\!\!\minuso\!\!\!\!-(W_{1},X)-\!\!\!\!\minuso\!\!\!\!-(W_{2},Y,Z)$ will hold. Thus,
$\Phi(X,W_{1};Y,W_{2}\|Z)\leq S(X^{\prime};W2,Y\|Z)+I(W_{1},X;W_{2},Y,Z\|X^{\prime})=S(X^{\prime};Y\|Z)+I(X;Y,Z\|X^{\prime})\leq\Phi(X;Y|Z)+\epsilon.$

### v3

@@ Appendix C An equivalent form of Conjecture 1
Note that by setting $Z=Z^{\prime}$, $Y=Y^{\prime}$, and replacing $X$ by $(X,X^{\prime})$ in Conjecture 1, we obtain the following equation (as a special case):
$S(X,X^{\prime};Y^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime},Z^{\prime};X|X^{\prime}),\qquad\forall P_{X,X^{\prime},Y^{\prime},Z^{\prime}}.$ (94)
Observe that $S(X,X^{\prime};Y^{\prime}\|Z^{\prime})\geq S(X,X^{\prime};X,Y^{\prime}\|X,Z^{\prime})$ because the first terminal can publicly reveal $X$ to all parties. Therefore, (94) can be weakened to
$S(X,X^{\prime};X,Y^{\prime}\|X,Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime},Z^{\prime};X|X^{\prime}),\qquad\forall P_{X,X^{\prime},Y^{\prime},Z^{\prime}}.$ (95)
@@
$S(X^{\prime};Y,Y^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X^{\prime},Z^{\prime};Y|Y^{\prime}).$ (96)
@@
$S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})-S(X,X^{\prime};Y^{\prime}|Z^{\prime})\leq I(X,X^{\prime},Z^{\prime};Y|Y^{\prime}).$ (97)
Adding (94) and (97), we obtain
$S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})\leq I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$ (98)
@@
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}|Z^{\prime})$ $=S(X;Y\|Z)-S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})+S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S(X;Y\|Z)-S(X;Y|Z^{\prime})+S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S_{\text{ow}}(X,Y\rightarrow Z^{\prime}|Z)+S(X,X^{\prime};Y,Y^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$ (99)
$\leq S_{\text{ow}}(X,Y\rightarrow Z^{\prime}|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime})$ (100)
@@
$\frac{1}{n}H(K_{B})$ $=\frac{1}{n}I(X^{n},F_{1:r},W_{1};K_{B})+\epsilon_{n}$ (102)
$=\frac{1}{n}(I(X^{n},F_{1:r},W_{1};K_{B})-I(K_{B};Z^{n},F_{1:r}))+2\epsilon_{n}$ (103)
$\leq\frac{1}{n}\Phi(X^{n},F_{1:r},W_{1};K_{B}\|Z^{n},F_{1:r})+2\epsilon_{n}$ (104)
$\leq\frac{1}{n}\Phi(X^{n},F_{1:r},W_{1};Y^{n},F_{1:r},W_{2}\|Z^{n},F_{1:r})+2\epsilon_{n}$ (105)
$\leq\frac{1}{n}\Phi(X^{n},F_{1:r-1},W_{1};Y^{n},F_{1:r-1},W_{2}\|Z^{n},F_{1:r-1})+2\epsilon_{n}$ (106)
$\leq\frac{1}{n}\Phi(X^{n},F_{1:r-2},W_{1};Y^{n},F_{1:r-2},W_{2}\|Z^{n},F_{1:r-2})+2\epsilon_{n}$ (107)
$\leq\cdots$ (108)
$\leq\frac{1}{n}\Phi(X^{n},W_{1};Y^{n},W_{2}\|Z^{n})+2\epsilon_{n}$ (109)
@@
where $\epsilon_{n}$ converges to zero as $\delta$ converges to zero, (102) follows from the fact that $K_{A}=K_{B}$ with probability $1-\delta$ and $H(K_{A}|X^{n},F_{1:r},W_{1})=0$, (103) follows from the privacy constraint, (104) from property iv in Lemma 4, (105) follows from property iii in Lemma 4 and the fact that $K_{B}$ is a function of $(Y^{n},W_{2},F_{1:r})$, (106)–(109) follow from property ii in Lemma 4 and the fact that $H(F_{i}|X^{n},F^{i-1},W_{1})=0$ for odd $i$ and $H(F_{i}|Y^{n},F^{i-1},W_{2})=0$ for even $i$, (110) follows from property v in Lemma 4, and (111) follows from property i in Lemma 4.
@@
Observe that $S(X;Y\|Z)$ satisfies similar properties as above, i.e., $S(X^{n};Y^{n}\|Z^{n})=nS(X;Y\|Z)$, and $S(F,X;F,Y\|F,Z)\leq S(F,X;Y\|Z)$, etc.; see [11].
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Proof of Property iv) We know that $S(X^{\prime};Y\|Z)\geq I(X^{\prime};Y)-I(Y;Z)$. Thus, $S(X^{\prime};Y\|Z)+I(X;Y,Z|X^{\prime})\geq I(X^{\prime};Y)-I(Y;Z)+I(X;Y,Z|X^{\prime})=I(X;Y)-I(Y;Z)+I(X;Z|X^{\prime},Y)\geq I(X;Y)-I(Y;Z).$ This completes the proof.
Proof of Property v) Take some arbitrary $P_{X^{\prime}|X}$ reaching within $\epsilon$ of the infimum defining $\Phi(X;Y|Z)$. If we define $(W_{1},W_{2})$ independent of $(X^{\prime},X,Y,Z)$, the Markov chain $X^{\prime}-\!\!\!\!\minuso\!\!\!\!-(W_{1},X)-\!\!\!\!\minuso\!\!\!\!-(W_{2},Y,Z)$ will hold. Thus,
$\Phi(X,W_{1};Y,W_{2}\|Z)\leq S(X^{\prime};W_{2},Y\|Z)+I(W_{1},X;W_{2},Y,Z|X^{\prime})=S(X^{\prime};Y\|Z)+I(X;Y,Z|X^{\prime})\leq\Phi(X;Y|Z)+\epsilon.$